OpenStudy (anonymous):

Find the exact value of the expression sin90 + cot45

OpenStudy (anonymous):

should we use the compound angle formulas?

OpenStudy (anonymous):

it says select the correct choice below, a) sin90+cot45= and b) undefined.

OpenStudy (anonymous):

thats what i am working with.

OpenStudy (anonymous):

whats your answer?

OpenStudy (anonymous):

what is sin 90?

Directrix (directrix):

sin90 + cot45 = 2 1 + 1 = 2

OpenStudy (anonymous):

so basically i need ot refer to my unit circle to get the ordered pairs. like for example, on cot of 45 the ordered pair \[\sqrt{2}/2,\sqrt{2}/2\] is 1 since it is divided?